### v2

@@ 1 Introduction
The difficulties of video collection and labeling complicates traditional data-driven training based on fully labeled datasets. Fortunately, few-shot action recognition (FSAR) improves learning efficiency and reduces the labeling dependency by classifying unseen actions from extremely few video samples. Therefore, FSAR has diverse real-world applications, including health monitoring and motion analysis (yan2023feature; wang2023openoccupancy). However, recognizing similar actions under regular viewpoint is a non-trival problem in FSAR. For instance, distinguishing “indoor climbing” and “construction working” is challenging, as subjects exhibit similar actions against a wall. To mitigate this issue, wide-angle videos provide contextual background, such as a “climbing wall” or a “construction site”, expressing actions within specific scenarios more accurately. According to established definitions (lai2021correcting; zhang2025madcow), wide-angle videos with a greater field of view (FoV) are widespread [^1] This work adopts the widely accepted definition of wide-angle as FoV exceeding 80.. FoV estimation (lee2021ctrl; hold2023perceptual) on popular FSAR benchmarks further reveals that approximately 35% of samples per dataset fall into this category, yet remain unexplored.
@@ 4.1 Experimental Configuration
According to TSN (wang2016temporal), each frame are sized into $3\times 256\times 256$ while $F$ of successive frames is set to 8. $3\times 224\times 224$ random crops and horizontal flipping data augmentation is applied during training while only the center crop is utilized in testing. As an exception, horizontal flipping is absent in SSv2 because of many actions with horizontal direction such as “Pulling S from left to right222“S” means “something”.”.
@@ Acknowledgments
The authors would like to appreciate all participants of peer review and cloud servers provided by Paratera Ltd. Wenbo Huang sincerely thanks those who offered companionship and encouragement during the most challenging times, even though life has since taken everyone on different paths. This work is supported by Science and Technology Major Special Program of Jiangsu Province under Grant No. BG2024028; Jiangsu Provincial Frontier Technology Research and Development Program under Grant No. BF2024070; National Natural Science Foundation of China under Grant Nos. 62472094, 62572119, 62232004, 62373194; Shenzhen Science and Technology Program under Grant No. KJZD20240903100814018; Jiangsu Provincial Key Laboratory of Network and Information Security under Grant No. BM2003201; Key Laboratory of Computer Network and Information Integration (Ministry of Education, China) under Grant No. 93K-9; Collaborative Innovation Center of Novel Software Technology and Industrialization; Collaborative Innovation Center of Wireless Communications Technology; the Fundamental Research Funds for the Central Universities; JSPS KAKENHI Nos. JP23K21676, JP24K02942, JP24K23849, JP25K21218, JP23K24851; JST PRESTO Grant No. JPMJPR23P5; JST CREST Grant No. JPMJCR21M2; JST NEXUS Grant No. JPMJNX25C4; and Startup Funds from The University of Newcastle, Australia.

### v3

@@ 1 Introduction
The difficulties of video collection and labeling complicates traditional data-driven training based on fully labeled datasets. Fortunately, few-shot action recognition (FSAR) improves learning efficiency and reduces the labeling dependency by classifying unseen actions from extremely few video samples. Therefore, FSAR has diverse real-world applications, including health monitoring and motion analysis (yan2023feature; wang2023openoccupancy). However, recognizing similar actions under regular viewpoint is a non-trivial problem in FSAR. For instance, distinguishing “indoor climbing” and “construction working” is challenging, as subjects exhibit similar actions against a wall. To mitigate this issue, wide-angle videos provide contextual background, such as a “climbing wall” or a “construction site”, expressing actions within specific scenarios more accurately. According to established definitions (lai2021correcting; zhang2025madcow), wide-angle videos with a greater field of view (FoV) are widespread***This work adopts the widely accepted definition of wide-angle as FoV exceeding 80.. FoV estimation (lee2021ctrl; hold2023perceptual) on popular FSAR benchmarks further reveals that approximately 35% of samples per dataset fall into this category, yet remain unexplored.
@@ 4.1 Experimental Configuration
According to TSN (wang2016temporal), each frame are sized into $3\times 256\times 256$ while $F$ of successive frames is set to 8. $3\times 224\times 224$ random crops and horizontal flipping data augmentation is applied during training while only the center crop is utilized in testing. As an exception, horizontal flipping is absent in SSv2 because of many actions with horizontal direction such as “Pulling S from left to right“S” means “something”.”.
@@ Acknowledgments
The authors would like to appreciate all participants of peer review and cloud servers provided by Paratera Ltd. Wenbo Huang sincerely thanks those who offered companionship and encouragement during the most challenging times, even though life has since taken everyone on different paths. This work is supported by Frontier Technologies Research and Development Program of Jiangsu under Grant No. BF2024070; National Natural Science Foundation of China under Grants Nos. 62472094, 62072099, 62232004, 62373194, 62276063; Jiangsu Provincial Key Laboratory of Network and Information Security under Grant No. BM2003201; Key Laboratory of Computer Network and Information Integration (Ministry of Education, China) under Grant No. 93K-9; the Fundamental Research Funds for the Central Universities; JSPS KAKENHI Nos. JP23K21676, JP24K02942, JP24K23849, JP25K21218, JP23K24851; JST PRESTO Grant No. JPMJPR23P5; JST CREST Grant No. JPMJCR21M2; JST NEXUS Grant No. JPMJNX25C4; and Startup Funds from The University of Newcastle, Australia.
Codehttps://github.com/wenbohuang1002/Otter